What is the magnitude of the electric field 0.3 m from a 10-4 C point charge?
1
Expert's answer
2012-12-05T10:03:45-0500
the magnitude is
E=kq/r^2 = 9*10^9 N*m^(2) C^(-2) *& 10^(-4)& C / (0.3 m)^2& = 10^7 V/m